We pour a 10l mixture of He, CO2 and CH4 in a container with NaOH. The volume of the gas that comes out is 7,24l. This gas that comes out reacts with 0.4 mol of Cl in presence of UV rays to give CCl4. Write the chemical reactions and determine the percentage of each ingredient in the 10l mixture.
1
Expert's answer
2020-09-21T06:49:21-0400

Out of these three gases, only CO2 reacts with NaOH:

2NaOH + CO2 --> Na2CO3 + H2O

Therefore, the volume and percentage of CO2 can be determined directly:

V(CO2) = 10L - 7.24L = 2.76 L

%CO2 = (2.76L / 10L) * 100% = 27.6%

The gas that reacts with Cl2 (there must be a mistake in the task description as chlorine exists only as a diatomic molecule) is CH4.

CH4 + 4Cl2 --> CCl4 + 4HCl

According to the reaction, 4 moles of Cl2 react with 1 mole of CH4. Therefore, 0.4 moles of Cl2 would react with 0.1 mole of CH4. Now the volume and percentage of CH4 can be determined (assuming STP conditions):

V (CH4) = n * VM = 0.1mol * 22.4L/mol = 2.24 L

%CH4 = (2.24L / 10L) * 100% = 22.4%

Finally, the remaining gas is He:

%He = 100% - 27.6% - 22.4% = 50%

Answer:

%CO2 = 27.6%

%CH4 = 22.4%

%He = 50%